# Duals and time-reversed flows of generalized Ornstein-Uhlenbeck processes

### v3

@@ 3.1 Siegmund duality on $\mathbb{R}$
Remark 3.6.
Let $(V_{t}^{x})_{t\geq 0}$ be the GOU process defined by (2.1), where $\Delta U_{t}\neq-1$ for all $t$. Then the condition $\Delta U_{t}>-1$ is not only sufficient for a dual process to $(V_{t}^{x})_{t\geq 0}$ in the sense of Definition 3.1 to exist as seen in the previous theorem, but also necessary. To see this, assume that $U$ has jumps of size less than $-1$ with positive probability. Then from (2.2), for $t>0$,
$\mathbb{P}(V_{t}^{x}\geq y)$ $=\mathbb{P}\bigg{(}\mathcal{E}(U)_{t}\left(x+\int_{(0,t]}\mathcal{E}(U)_{s-}^{-1}\mathop{}\!\mathrm{d}\eta_{s}\right)\geq y,\mathcal{E}(U)_{t}>0\bigg{)}$
$\quad+\mathbb{P}\bigg{(}\mathcal{E}(U)_{t}\left(x+\int_{(0,t]}\mathcal{E}(U)_{s-}^{-1}\mathop{}\!\mathrm{d}\eta_{s}\right)\geq y,\mathcal{E}(U)_{t}<0\bigg{)}$
which converges to $\mathbb{P}(\mathcal{E}(U)_{t}>0)$ as $x\to\infty$, and the latter probability is in $(0,1)$ by (2.3). However, if a dual process $(R_{t}^{y})_{t\geq 0}$ as in Definition 3.1 were to exist, then $\lim_{x\to\infty}\mathbb{P}(V_{t}^{x}\geq y)=1$ for all $y\in\mathbb{R}$ by (3.1), since we only allowed $\mathbb{R}$-valued dual processes. Even if we had allowed [,][$\infty$,$\infty$][ - ∞ , ∞ ]-valued dual processes $(R_{t}^{y})_{t\geq 0}$ in Definition 3.1, then still a dual Markov process $(R_{t}^{y})_{t\geq 0}$ would not exist. For if there were such an [,][$\infty$,$\infty$][ - ∞ , ∞ ]-valued Markov process $(R_{t}^{y})$, then $\mathbb{P}(R_{t}^{y}<\infty)=\mathbb{P}(\mathcal{E}(U)_{t}>0)$ for $y\in\mathbb{R}$ by the same argument, hence
$\mathbb{P}(\mathcal{E}(U)_{t}>0)$ $\geq\mathbb{P}(R_{t}^{y}\leq x)=\mathbb{P}(V_{t}^{x}\geq y)=\mathbb{P}\bigg{(}\mathcal{E}(U)_{t}\left(x+\int_{(0,t]}\mathcal{E}(U)_{s-}^{-1}\,\mathop{}\!\mathrm{d}\eta_{s}\right)\geq y\bigg{)}$
for all $x,y\in\mathbb{R}$, and the latter expression converges to $1$ as $y\to-\infty$, a contradiction.   $\Box$
3.2 On the relation between dual GOU processes and hitting times
As described in the introduction, Siegmund’s duality is often used to match ruin probabilities, i.e. hitting probabilities of the negative half line, with stationary distributions of the dual processes. For the GOU process Paulsen [16, Thm. 3.2] provides an expression for its ruin probability in the case that the driving processes have finite jump activity. His result has been generalized to the setting of this paper in [3, Thm. 2.4] where it reads as follows.
@@ 3.2 On the relation between dual GOU processes and hitting times
Consider the GOU process $(V_{t}^{x})_{t\geq 0}$ solving (2.1) for some bivariate Lévy process $(U,L)$ with $\Delta U_{t}>-1$ for all $t$, where $x>0$. Assume that $|\mathcal{E}(U)_{t}|$ converges almost surely to +$\infty$+ ∞ as $t\to\infty$ and that the integral $\int_{(0,t]}\mathcal{E}(U)_{s-}^{-1}\,\mathrm{d}\eta_{s}$ converges almost surely as $t\to\infty$ to a non-degenerate random variable $V_{\infty}=\int_{0}^{\infty}\mathcal{E}(U)_{s-}^{-1}\,\mathrm{d}\eta_{s}$ with distribution function $H$. Let $\tau(x)=\inf\{t\geq 0,V_{t}^{x}\leq 0\}$. Then
$\mathbb{P}(\tau(x)<\infty)\,\mathbb{E}[H(-V_{\tau(x)}^{x})|\tau(x)<\infty]=H(-x),$
where $\mathbb{E}[\cdot|\tau(x)<\infty]$ is interpreted as 00 if $\mathbb{P}(\tau(x)<\infty)=0$.
Under the conditions and using the same notations as in Theorem 3.5 and Proposition 3.7, we note that
$H(-x)=\mathbb{P}\left(\int_{0}^{\infty}\mathcal{E}(U)_{s-}^{-1}\,\mathrm{d}\eta_{s}\leq-x\right)=\mathbb{P}\left(\int_{0}^{\infty}\mathcal{E}(U)_{s-}^{-1}\,\mathrm{d}K_{s}\geq x\right).$
As mentioned in the Preliminaries, by [6, Thm. 2.1], in case of its convergence the law of the exponential functional
$\int_{0}^{\infty}\mathcal{E}(U)_{s-}^{-1}\,\mathrm{d}K_{s}=-V_{\infty},$
is the unique non-causal stationary distribution of $(V_{t}^{x})_{t\geq 0}$, as well as the unique causal stationary distribution of the dual GOU process $(R_{t}^{y})_{t\geq 0}$. Moreover, $R_{t}^{y}$ converges in distribution to $R_{\infty}:=-V_{\infty}$ as $t\to\infty$ for every $y\in\mathbb{R}$. This suggests to interpret duals of GOU processes as a sort of time-reversal and motivates our study of time-reversed GOU processes in the Section 4. Further, these observations lead to the following corollary.
Corollary 3.8.
Under the conditions of Proposition 3.7 the limit of $\mathbb{P}(V_{t}^{x}\leq 0)$ as $t\to\infty$ exists and is equal to $H(-x)=\mathbb{P}(R_{\infty}\geq x)$, i.e.
$\lim_{t\to\infty}\mathbb{P}(V_{t}^{x}\leq 0)=\mathbb{P}(R_{\infty}\geq x)=\mathbb{P}(\tau(x)<\infty)\,\mathbb{E}[H(-V_{\tau(x)}^{x})|\tau(x)<\infty].$
In particular, if $-L$ is a subordinator, then
$\mathbb{P}(\tau(x)<\infty)=\mathbb{P}(R_{\infty}\geq x)=H(-x).$
Proof.
By Theorem 3.5, not only is $(R_{t}^{y})_{t\geq 0}$ dual to $(V_{t}^{x})_{t\geq 0}$, but also $(V_{t}^{x})_{t\geq 0}$ is dual to $(R_{t}^{y})_{t\geq 0}$. Hence $\mathbb{P}(V_{t}^{x}\leq 0)=\mathbb{P}(R_{t}^{0}\geq x)$ and the latter converges by assumption as $t\to\infty$ to $\mathbb{P}(R_{\infty}\geq x)=\mathbb{P}(-V_{\infty}\geq x)=H(-x)$ (recall that $H$ is continuous). The formula for the limit is then immediate from Proposition 3.7.
For the second statement it suffices to assume that $\mathbb{P}(\tau(x)<\infty)>0$. Note that if $-L$ is a subordinator, also $-\eta$ is a subordinator by the observation just below (2.4), which implies that $\int_{0}^{\infty}\mathcal{E}(U)_{s-}^{-1}\,\mathrm{d}\eta_{s}\leq 0$ a.s. and hence $H(0)=1$. Since $V_{\tau(x)}^{x}\leq 0$ on $\{\tau(x)<\infty\}$ by the càdlàg paths of $V^{x}$ this implies $H(-V_{\tau(x)}^{x})=1$ and hence the statement by Proposition 3.7. ∎
3.3 Non-negative Siegmund dual processes
As mentioned before, if $\Delta U_{t}>-1$ for all $t\geq 0$ and $L$ is a subordinator, then $(V_{t}^{x})_{t\geq 0}$ is a time-homogeneous Markov process with state-space $[0,\infty)$, if one restricts to non-negative starting values $x\in[0,\infty)$. Thus, according to Siegmund’s original work, this GOU process has a dual process on $[0,\infty]$ in the sense of (1.1) that will be derived in the next proposition.
Proposition 3.9.
Let $(V_{t}^{x})_{t\geq 0}$ with $x\geq 0$ be a solution of the SDE (2.1) for the bivariate Lévy process $(U_{t},L_{t})_{t\geq 0}$ with $\Delta U_{t}>-1$ for all $t$ and $L$ being a subordinator. Let $(R_{t}^{y})_{t\geq 0}$ with $y\geq 0$ be the solution of (3.3) and set
@@ 3.3 Non-negative Siegmund dual processes
$\widehat{R}_{t}^{y}:=R_{t}^{y}\mathds{1}_{t<\tau_{R}(y)},$
is a time-homogeneous Markov process on $[0,\infty)$, and it is dual to $(V_{t}^{x})_{t\geq 0}$, $x\geq 0$, in the sense of (1.1). Moreover, we have $\widehat{R}_{t}^{y}=\max\{R_{t}^{y},0\}$ for $t,y\geq 0$.
@@
and that $R_{t}^{y}$ is explicitly given by (3.8). As $(R_{t}^{y})_{t\geq 0}$ being a GOU process - is a time-homogeneous Markov process, its killed version $(\widehat{R}_{t}^{y})_{t\geq 0}$ is a time-homogeneous Markov process as well. Moreover, under the given conditions, $\mathcal{E}(W)_{t}=\mathcal{E}(U)_{t}^{-1}>0$ for all $t\geq 0$, and as $L$ is a subordinator it follows from (3.8) that
@@
Due to time-homogeneity of the GOU process $(R_{t}^{y})_{t\geq 0}$ this implies for any $y\in\mathbb{R}$ that $R_{t}^{y}\leq 0$ for all $t\geq\tau_{R}(y)$, and in particular
$\widehat{R}_{t}^{y}=R_{t}^{y}\mathds{1}_{t<\tau_{R}(y)}=\max\{R_{t}^{y},0\},\quad t\geq 0.$
@@
provided $\eta$ is not the zero-subordinator. Since the function $[0,\infty)\to[0,1]$, $y\mapsto\mathbb{P}(\widehat{R}_{t}^{y}\geq x)$, is nondecreasing and right-continuous for any $t,x\geq 0$, by Siegmund’s result the process $(\widehat{R}_{t}^{y})_{t\geq 0}$, $y\geq 0$, has a dual on $[0,\infty]$, but it is not $(V_{t}^{x})_{t\geq 0}$, $x\geq 0$.
4 The time-reversed flow of the generalized Ornstein-Uhlenbeck process
@@
In contrast to duality which can be considered as a distributional concept, time-reversals of stochastic processes are typically defined pathwise. Therefore we use a different framework in this section than before, and study the GOU process in terms of its stochastic flow and the reversed stochastic flow. Recall the definition of the time-reversal $\widetilde{X}$ of some càdlàg process $X=(X_{s})_{s\in[0,t]}$ as
$\widetilde{X}_{s}=\begin{cases}0,&\text{if }s=0,\\ X_{(t-s)-}-X_{t-},&\text{if }0<s<t,\\ X_{0}-X_{t-},&\text{if }s=t.\end{cases}$ (4.1)
@@
For $t=1$ this coincides with the definition given in [18, Chapter VI.4]. It is well known that when $X$ is a Lévy process, then $(\widetilde{X}_{s})_{s\in[0,t]}$ is also a Lévy process, with the same distribution as $(-X_{s})_{s\in[0,t]}$, see e.g. [7, Lem. II.2], [8, Thm. 11.4] or [19, Prop. 41.8]. Stochastic integrals with respect to time-reversed Lévy processes are then to be seen with respect to their augmented natural filtration, for which the time-reversed Lévy processes are semimartingales.
@@
As before, let $(V_{t}^{x})_{t\geq 0}$ be a solution of the SDE (2.1) for the bivariate Lévy process $(U_{t},L_{t})_{t\geq 0}$. As no stochastic monotonicity is needed in the context of time-reversals, we shall only assume that $\Delta U_{t}\neq-1$ for all $t$.
@@
Let $(V_{s}^{x})_{s\geq 0}$ be a solution of the SDE (2.1) for the bivariate Lévy process $(U_{s},L_{s})_{s\geq 0}$ with $\Delta U_{s}\neq-1$ for all $s\geq 0$, as given in (2.2) with $\eta$ defined in (2.4), where $x\in\mathbb{R}$. Let $t>0$ be fixed. Denote the time-reversal of $(U,L,\eta)$ on $[0,t]$ by $(\widetilde{U},\widetilde{L},\widetilde{\eta})$, as defined in (4.1) and define
@@
Then $(T_{s},\widetilde{L}_{s},\widetilde{\eta}_{s})_{s\in[0,t]}$ is a Lévy process on $[0,t]$ and, almost surely, we have for all $s\in[0,t]$
@@
Further, the stochastic process corresponding to the reversed stochastic flow induced by (4.2), i.e. the process $(R_{s}^{y})_{s\in[0,t]}$, $y\in\mathbb{R}$, defined by
@@
is the GOU which is the unique solution of the SDE
@@
Since the right-hand sides of both (4.2) and (4.6) define càdlàg processes in $s$, it is enough to show that for given $s$ they agree almost surely (where the exceptional null set may depend on $s$). Since $\eta$ almost surely does not jump at the fixed time $t-s$, the integral over the compact interval $[t-s,s]$ in (4.6) is almost surely equal to the corresponding integral over the half-open interval $(t-s,s]$. Hence, by (4.6) and (4.7), Equation (4.2) will follow if we can show that, almost surely for each fixed $s\in[0,t]$,
$-\frac{\mathcal{E}(U)_{t}}{\mathcal{E}(U)_{t-s}}\int_{(t-s,t]}\left(\frac{\mathcal{E}(U)_{u-}}{\mathcal{E}(U)_{t-s}}\right)^{-1}\,\mathrm{d}\eta_{u}=\int_{(0,s]}\mathcal{E}(T)_{u-}^{-1}\mathop{}\!\mathrm{d}\widetilde{L}_{u}.$ (4.8)
For the treatment of (4.8), define for fixed $s\in[0,t]$
@@
the negative of the time reversal of $(U^{\prime},\eta^{\prime})$ at $s$. Then, with [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denoting quadratic covariation,
@@
this is Lemma 6.1 in [12] when all jumps of $U$ are greater than $-1$, and for general $U$ it follows from the proof of Proposition 8.3 in [4] (the left-hand side is the probability limit of the $-B^{\sigma}$ appearing in that proof, while the right-hand side is the probability limit of $A^{\sigma}$). Since $\mathcal{E}(\widehat{U^{\prime}})_{r}=1+\int_{(0,r]}\mathcal{E}(\widehat{U^{\prime}})_{u-}\mathop{}\!\mathrm{d}\widehat{U_{u}^{\prime}}$ we can rewrite the left-hand side of (4.8) as
@@
From  Theorems 3.5 and 4.1 we deduce that if all jumps of $U$ are greater than $-1$, then the process $(R_{s}^{y})_{s\in[0,t]}$, $y\in\mathbb{R}$, given in (4.4) is Sigmund-dual to $(V_{s}^{x})_{s\in[0,t]}$, $x\in\mathbb{R}$, with the obvious notion of Sigmund-duality for processes defined only on $[0,t]$.

### v4

@@ 3.1 Siegmund duality on $\mathbb{R}$
We can relate the stationary distribution of the GOU $(V_{t}^{x})_{t\geq 0}$ to that of its dual process $(R_{t}^{y})_{t\geq 0}$, provided they exist.
Corollary 3.6.
Suppose that (2.5) is satisfied, let $(V_{t}^{x})_{t\geq 0}$ be the GOU process given by (2.2), and $(R_{t}^{y})_{t\geq 0}$ its dual process as specified in (3.3).
  1. (a)
    Suppose $\lim_{t\to\infty}\mathcal{E}(U)_{t}=0$ almost surely and that $V_{\infty}:=\int_{0}^{\infty}\mathcal{E}(U)_{s-}\,\mathop{}\!\mathrm{d}L_{s}$ exists almost surely and is finite. Then the distribution of $V_{\infty}$ is both the stationary distribution of the causal GOU process $(V_{t}^{x})_{t\geq 0}$ as well as the stationary distribution of the non-causal GOU process $(R_{t}^{y})_{t\geq 0}$.
  2. (b)
    Suppose $\lim_{t\to\infty}\mathcal{E}(U)_{t}^{-1}=0$ almost surely and that $R_{\infty}:=\int_{0}^{\infty}\mathcal{E}(U)_{s-}^{-1}\,\mathop{}\!\mathrm{d}K_{s}$ exists almost surely and is finite. Then the distribution of $R_{\infty}$ is both the stationary distribution of the non-causal GOU process $(V_{t}^{x})_{t\geq 0}$ as well as the stationary distribution of the causal GOU process $(R_{t}^{y})_{t\geq 0}$.
Proof.
Recall (3.5) from which $\lim_{t\to\infty}\mathcal{E}(W)_{t}^{-1}=\lim_{t\to\infty}\mathcal{E}(U)_{t}$. Further, the right-hand side of (2.3) applied to $(W_{t},K_{t})_{t\geq 0}$ instead of $(U_{t},L_{t})_{t\geq 0}$ gives
$K_{t}-\sum_{0<s\leq t}\frac{\Delta W_{s}\Delta K_{s}}{1+\Delta W_{s}}-t\sigma_{W,K}=-L_{t},$
where we used the definition of $K_{t}$, and Equations (3.6) and (2.3), from which $\Delta W_{s}=\frac{-\Delta U_{s}}{1+\Delta U_{s}}$, $\Delta K_{s}=-\Delta\eta_{s}=\frac{-\Delta L_{s}}{1+\Delta U_{s}}$, and $\sigma_{W,K}=\sigma_{U,L}$. Assertion (a) is then an immediate consequence of Lemma 2.4. The proof of (b) is similar. ∎
As motivated in Example 2.6 we aim to investige the connection between the limiting distribution of $(R_{t}^{y})_{t\geq 0}$ and the ruin probability of $(V_{t}^{x})_{t\geq 0}$. This can now be done by combining Proposition 2.7 with our results.
Corollary 3.7.
Under the conditions and with the same notations as in Proposition 2.7, the limit of $\mathbb{P}(V_{t}^{x}\leq 0)$ as $t\to\infty$ exists and is equal to $H(-x)=\mathbb{P}(R_{\infty}\geq x)$, so that
$\lim_{t\to\infty}\mathbb{P}(V_{t}^{x}\leq 0)=\mathbb{P}(R_{\infty}\geq x)=\mathbb{P}(\tau(x)<\infty)\,\mathbb{E}[H(-V_{\tau(x)}^{x})|\tau(x)<\infty],\quad x\geq 0,$ (3.9)
where $R_{\infty}=\int_{0}^{\infty}\mathcal{E}(U)_{s-}^{-1}\,\mathrm{d}K_{s}$ as in Corollary 3.6 (b). In particular, if L-L is a subordinator, then
$\mathbb{P}(\tau(x)<\infty)=\mathbb{P}(R_{\infty}\geq x)=H(-x),\quad x\geq 0.$
Proof.
By Theorem 3.5, not only is $(R_{t}^{y})_{t\geq 0}$ dual to $(V_{t}^{x})_{t\geq 0}$, but also $(V_{t}^{x})_{t\geq 0}$ is dual to $(R_{t}^{y})_{t\geq 0}$. Hence $\mathbb{P}(V_{t}^{x}\leq 0)=\mathbb{P}(R_{t}^{0}\geq x)$. Since the distribution of $R_{\infty}=\int_{0}^{\infty}\mathcal{E}(U)_{s-}^{-1}\,\mathrm{d}K_{s}=-\int_{0}^{\infty}\mathcal{E}(U)_{s-}^{-1}\,\mathrm{d}\eta_{s}$ is the stationary distribution of a causal GOU process which by assumption is non-degenerate, it cannot have atoms by [8, Thm. 2.2]. Since $R_{t}^{0}$ converges in distribution to $R_{\infty}$ as $t\to\infty$ by Lemma 2.4 (a), we conclude that $\mathbb{P}(V_{t}^{x}\leq 0)$ converges to $\mathbb{P}(R_{\infty}\geq x)=\mathbb{P}(-R_{\infty}\leq-x)=H(-x)$. The formula (3.9) for the limit is then immediate from Proposition 2.7.
For the second statement it suffices to assume that $\mathbb{P}(\tau(x)<\infty)>0$, for if $\mathbb{P}(\tau(x)<\infty)=0$ then also $\mathbb{P}(R_{\infty}\geq x)=0$ by (3.9) and the claim follows. Note that if L-L is a subordinator, also $\eta$ is a subordinator by the observation just below (2.6), which implies that $\int_{0}^{\infty}\mathcal{E}(U)_{s-}^{-1}\,\mathrm{d}\eta_{s}\leq 0$ a.s. and hence $H(0)=1$. Since $V_{\tau(x)}^{x}\leq 0$ on $\{\tau(x)<\infty\}$ by the càdlàg paths of $(V^{x}_{t})_{t\geq 0}$ this implies $H(-V_{\tau(x)}^{x})=1$ and hence the statement when $\mathbb{P}(\tau(x)<\infty)>0$ by (3.9). ∎
3.2 Non-negative Siegmund dual processes
As mentioned before, if $\Delta U_{t}>-1$ for all $t\geq 0$ and $L$ is a subordinator, then $(V_{t}^{x})_{t\geq 0}$ is a time-homogeneous Markov process with state-space $0,\infty)$, if one restricts to non-negative starting values $x\in[0,\infty)$. Thus, according to Siegmund’s original work, this GOU process has a dual process on $0,\infty]$ in the sense of (1.1) that will be derived in the next proposition.
@@ 3.2 Non-negative Siegmund dual processes
Assume (2.5) and let $(V_{t}^{x})_{t\geq 0}$ with $x\geq 0$ be the GOU given by (2.2). Further assume that $L$ is a subordinator. Let $(R_{t}^{y})_{t\geq 0}$ with $y\geq 0$ be the solution of (3.3) and set
@@
$\widehat{R}_{t}^{y}:=R_{t}^{y}\mathds{1}_{\{t<\tau_{R}(y)\}},$
is a time-homogeneous Markov process on $0,\infty)$, and it is dual to $(V_{t}^{x})_{t\geq 0}$, $x\geq 0$, in the sense of (1.1). Moreover, we have $\widehat{R}_{t}^{y}=\max\{R_{t}^{y},0\}$ for $t,y\geq 0$.
@@
and that $R_{t}^{y}$ is explicitly given by (3.8). As $(R_{t}^{y})_{t\geq 0}$ – being a GOU process – is a time-homogeneous Markov process, its killed version $(\widehat{R}_{t}^{y})_{t\geq 0}$ is a time-homogeneous Markov process as well. Moreover, under the given conditions, $\mathcal{E}(W)_{t}=\mathcal{E}(U)_{t}^{-1}>0$ for all $t\geq 0$, and as $L$ is a subordinator it follows from (3.8) that
@@
Due to time-homogeneity of $(R_{t}^{y})_{t\geq 0}$ this implies $R_{t}^{y}\leq 0$ for any $y\in\mathbb{R}$, and for all $t\geq\tau_{R}(y)$. In particular
$\widehat{R}_{t}^{y}=R_{t}^{y}\mathds{1}_{\{t<\tau_{R}(y)\}}=\max\{R_{t}^{y},0\},\quad t\geq 0.$
@@
provided $\eta$ is not the zero-subordinator. Since the function $0,\infty)\to[0,1]$, $y\mapsto\mathbb{P}(\widehat{R}_{t}^{y}\geq x)$, is nondecreasing and right-continuous for any $t,x\geq 0$, by Siegmund’s result the process $(\widehat{R}_{t}^{y})_{t\geq 0}$, $y\geq 0$, has a dual on $0,\infty]$, but it is not $(V_{t}^{x})_{t\geq 0}$, but instead the process $(\widehat{V}_{t}^{x})_{t\geq 0}$ defined as
$\widehat{V}_{t}^{x}=V_{t}^{x}\mathds{1}_{\{t<\tau_{V}(x)\}},\quad\text{with }\tau_{V}(x):=\inf\{t\geq 0:V_{t}^{x}\leq 0\}.$
To see this, observe that $\widehat{V}_{t}^{0}\equiv 0$ which implies
$\mathbb{P}(\widehat{R}^{y}_{t}\geq 0)=1=\mathbb{P}(\widehat{V}_{t}^{0}\leq y)\quad\text{for all }y\geq 0,t\geq 0.$
Further, for $x>0$, as $L$ and hence $\eta$ is a subordinator, $V_{t}^{x}>0$ for all $t\geq 0$, implying $\tau_{V}(x)=\infty$ almost surely. With this it follows from Theorem 3.5 and Proposition 3.8 for all $y\geq 0,x>0$ and $t\geq 0$
$\mathbb{P}(\widehat{R}^{y}_{t}\geq x)$ $=\mathbb{P}(\max\{R_{t}^{y},0\}\geq x)=\mathbb{P}(R_{t}^{y}\geq 0)$
$=\mathbb{P}(V_{t}^{x}\leq y)=\mathbb{P}(V_{t}^{x}\mathds{1}_{\{t<\tau_{V}(x)\}}\leq y)=\mathbb{P}(\widehat{V}_{t}^{x}\leq y).$
Hence $(\widehat{V}_{t}^{x})_{t\geq 0}$ is dual to $(\widehat{R}_{t}^{y})_{t\geq 0}$ in the sense of (1.1), but again the converse statement fails as can easily be checked considering the duality relation for $x=0$.
As a consequence of Proposition 3.8 we can relate the ruin probability of $(R_{t}^{y})_{t\geq 0}$ to the limit distribution of $(V_{t}^{0})_{t\geq 0}$ in the subordinator case.
Corollary 3.10.
Assume (2.5) and let $(V_{t}^{x})_{t\geq 0}$, $(R_{t}^{y})_{t\geq 0}$ and $\tau_{R}(y)$ be as in Proposition 3.8 with $L$ being a subordinator. Then
$\mathbb{P}(\tau_{R}(y)<\infty)=\lim_{t\to\infty}\mathbb{P}(V_{t}^{0}\geq y)\quad\text{for every }y\geq 0.$
Denote $V_{\infty}:=\int_{0}^{\infty}\mathcal{E}(U)_{s}\,\mathop{}\!\mathrm{d}L_{s}$, which is a $0,\infty]-$valued numerical random variable (in case of almost sure convergence, it is finite). Then, if neither $U$ nor $L$ is the zero-process and if there exists no $k\in\mathbb{R}\setminus\{0\}$ such that (2.11) holds, then
$\mathbb{P}(\tau_{R}(y)<\infty)=\mathbb{P}(V_{\infty}\geq y)$
for every $y\geq 0$. Finally, if additionally $\lim_{t\to\infty}\mathcal{E}(U)_{t}=0$ almost surely, then
$\mathbb{P}(\tau_{R}(y)<\infty)=\lim_{t\to\infty}\mathbb{P}(V_{t}^{x}\geq y)$
for all $x,y\geq 0$.
Proof.
Observe that $(\widehat{R}_{t}^{y})$ as defined in Proposition 3.8 stays at 0 once it has reached 0. Since $(\widehat{R}_{t}^{y})_{t\geq 0}$ is dual to $(V_{t}^{x})_{t\geq 0}$ in the sense of (1.1), for fixed $y\geq 0$ we have
$\mathbb{P}(\tau_{R}(y)<\infty)$ $=\mathbb{P}(\exists\,t\geq 0:\widehat{R}_{t}^{y}=0)=\lim_{t\to\infty}\mathbb{P}(\widehat{R}_{t}^{y}\leq 0)=\lim_{t\to\infty}\mathbb{P}(V_{t}^{0}\geq y).$
Let $\xi_{t}=-\log\mathcal{E}(U)_{t}$ as in (2.6), then according to [19, Thm. 2.1], $\int_{(0,t]}\mathrm{e}^{-\xi_{s-}}\mathop{}\!\mathrm{d}L_{s}=\int_{(0,t]}\mathcal{E}(U)_{s-}^{-1}\,\mathrm{d}L_{s}$ either converges a.s. to a finite random variable as $t\to\infty$, or diverges to $\infty$ in probability (and hence almost surely since $L$ is a subordinator), or (2.12) holds for some $k\neq 0$. But (2.12) is equivalent to (2.11) which has been ruled out by assumption. Hence $V_{\infty}$ is either finite almost surely or almost surely equal to $\infty$. When $V_{\infty}$ is finite almost surely, its distribution cannot have atoms by [8, Thm. 2.2] since (2.11) has been ruled out. Hence, in all cases the distribution of $V_{\infty}$ has no atoms on the real line.
From (2.2) and Lemma 2.2 we see that $V_{t}^{x}$ has the same distribution as $\mathcal{E}(U)_{t}x+\int_{(0,t]}\mathcal{E}(U)_{s-}\,\mathop{}\!\mathrm{d}L_{s}$, which converges almost surely to $V_{\infty}$ as $t\to\infty$ if $x=0$, or if $x>0$ and additionally $\mathcal{E}(U)_{t}$ converges to 0 as $t\to\infty$. Since $V_{\infty}$ has no atom at $y$, we then have $\lim_{t\to\infty}\mathbb{P}(V_{t}^{x}\geq y)=\mathbb{P}(V_{\infty}\geq y)$ in both cases, giving the claim. ∎
4 The inverse flow of the generalized Ornstein-Uhlenbeck process
@@
In contrast to duality, which can be considered a distributional concept, time-reversals of stochastic processes are typically defined pathwise. Therefore, we use a different framework in this section than before and study the GOU process in terms of its stochastic flow and the inverse stochastic flow. This is possible because, in the GOU setting, the structure of the SDE admits an explicit inverse mapping for the jumps, rather than applying the generalized inverse used in [2]. In comparison with the time-reversed SDE that can be derived via [43, Thm. VI.4.22], in Theorem 4.1 we obtain the same SDE for the inverse stochastic flow. Similar considerations appear at the end of Chapter VI.4 in [43, Proof of Thm. VI.4.23], where time-reversals of SDEs of Itô diffusions are considered.
Recall the definition of the time-reversal $\widetilde{X}$ at $t>0$ of some càdlàg process $X=(X_{s})_{s\in[0,t]}$ as
[…]
@@
For $t=1$ this coincides with the definition given in [43, Chapter VI.4]. It is well known that when $X$ is a Lévy process, then $(\widetilde{X}_{s})_{s\in[0,t]}$ is also a Lévy process, with the same distribution as $(-X_{s})_{s\in[0,t]}$, see e.g. [9, Lem. II.2], [11, Thm. 11.4] or [44, Prop. 41.8]. Stochastic integrals with respect to time-reversed Lévy processes are then to be seen with respect to their augmented natural filtration, for which the time-reversed Lévy processes are semimartingales.
@@
As before, let $(V_{t}^{x})_{t\geq 0}$ be a solution of the SDE (1.2) for the bivariate Lévy process $(U_{t},L_{t})_{t\geq 0}$. As no stochastic monotonicity is needed in the context of time-reversals, we shall only assume (2.1), i.e. $\Delta U_{t}\neq-1$ for all $t$.
@@
[…]
@@
Then $(T_{s},\widetilde{L}_{s},\widetilde{\eta}_{s})_{s\in[0,t]}$ is a Lévy process on $[0,t]$ and, almost surely, for all $s\in[0,t]$
@@
Further, the stochastic process corresponding to the inverse stochastic flow induced by (4.2), i.e. the process $(R_{s}^{y})_{s\in[0,t]}$, $y\in\mathbb{R}$, defined by
@@
is the GOU process which is the unique solution of the SDE
@@
Since the right-hand sides of both (4.2) and (4.6) define càdlàg processes in $s$, it is enough to show that for given $s$ they agree almost surely (where the exceptional null set may depend on $s$). Since $\eta$ almost surely does not jump at the fixed time tst-s, the integral over the compact interval [ts,s][t-s,s] in (4.6) is almost surely equal to the corresponding integral over the half-open interval (ts,s](t-s,s]. Hence, by (4.6) and (4.7), Equation (4.2) will follow if we can show that, almost surely for each fixed $s\in[0,t]$,
$\frac{\mathcal{E}(U)_{t}}{\mathcal{E}(U)_{t-s}}\int_{(t-s,t]}\left(\frac{\mathcal{E}(U)_{u-}}{\mathcal{E}(U)_{t-s}}\right)^{-1}\,\mathrm{d}\eta_{u}=\int_{(0,s]}\mathcal{E}(T)_{u-}^{-1}\mathop{}\!\mathrm{d}\widetilde{L}_{u}.$ (4.8)
For this, define for fixed $s\in[0,t]$
@@
the negative of the time-reversal of $(U^{\prime},\eta^{\prime})$ at $s$. Then, with $\cdot$$\cdot$] denoting quadratic covariation,
@@
this is Lemma 6.1 in [34] when all jumps of $U$ are greater than 1-1, and for general $U$ it follows from the proof of Proposition 8.3 in [6] (the left-hand side is the probability limit of the $B^{\sigma}$ appearing in that proof, while the right-hand side is the probability limit of $A^{\sigma}$). Since $\mathcal{E}(\widehat{U^{\prime}})_{r}=1+\int_{(0,r]}\mathcal{E}(\widehat{U^{\prime}})_{u-}\mathop{}\!\mathrm{d}\widehat{U_{u}^{\prime}}$ we can rewrite the left-hand side of (4.8) as
@@
From Theorems 3.5 and 4.1 we deduce that, if all jumps of $U$ are greater than 1-1, then the process $(R_{s}^{y})_{s\in[0,t]}$, $y\in\mathbb{R}$, given in (4.4) is Siegmund dual to $(V_{s}^{x})_{s\in[0,t]}$, $x\in\mathbb{R}$, with the obvious notion of Siegmund duality for processes defined only on $[0,t]$.